Old Country Links, Inc., produces sausages in three production departments—Mixing, Casing and Curing, and Packaging. In the Mixing Department, meats are prepared and ground and then mixed with spices. The spiced meat mixture is then transferred to the Casing and Curing Department, where the mixture is force-fed into casings and then hung and cured in climate-controlled smoking chambers. In the Packaging Department, the cured sausages are sorted, packed, and labeled. The company uses the weighted-average method in its process costing system. Data for September for the Casing and Curing Department follow:

Percent Completed
Units Mixing Materials Conversion
Work in process inventory, September 1 7 100 % 60 % 50 %
Work in process inventory, September 30 7 100 % 20 % 10 %


Mixing Materials Conversion
Work in process inventory, September 1 $ 13,692 $ 35 $ 3,563
Cost added during September $ 205,548 $ 19,947 $ 163,181


Mixing cost represents the costs of the spiced meat mixture transferred in from the Mixing Department. The spiced meat mixture is processed in the Casing and Curing Department in batches; each unit in the above table is a batch and one batch of spiced meat mixture produces a set amount of sausages that are passed on to the Packaging Department. During September, 109 batches (i.e., units) were completed and transferred to the Packaging Department.

Required:

1. Determine the Casing and Curing Department's equivalent units of production for mixing, materials, and conversion for the month of September.

2. Compute the Casing and Curing Department's cost per equivalent unit for mixing, materials, and conversion for the month of September.

3. Compute the Casing and Curing Department's cost of ending work in process inventory for mixing, materials, conversion, and in total for September.

4. Compute the Casing and Curing Department's cost of units transferred out to the Packaging Department for mixing, materials, conversion, and in total for September.

5. Prepare a cost reconciliation report for the Casing and Curing Department for September.
Answer #1

1) Equivalent unit

Mixing Material Conversion
Unit transferred out 109 109 109
Ending WIP 7 7*20% = 1.40 7*10% = 0.70
Equivalent unit 116 110.40 109.70

2) Cost per equivalent unit

Mixing Material Conversion
Beginning WIP 13692 35 3563
Cost added 205548 19947 163181
Total cost 219240 19982 166744
Equivalent unit 116 110.40 109.70
Cost per equivalent unit 1890 181 1520

3) Cost of ending WIP

Mixing Material Conversion Total
Cost of ending WIP 7*1890 = 13230 1.4*181 = 253 0.7*1520 = 1064 14547

4) Cost of unit transferred out

Mixing Material Conversion Total
Cost of unit transferred out 109*1890 = 206010 19729 165680 391419

5) Cost reconciliation

Cost accounted for
Beginning WIP 17290
Cost added 388676
Total cost accounted for 405966
Cost accounted as
Cost of unit transferred out 391419
Cost of ending WIP 14547
Total cost accounted as 405966